OpenStudy (anonymous):

5x+10=10y what is this equall

OpenStudy (anonymous):

x+2=2y

OpenStudy (anonymous):

y= 1/2x +1

OpenStudy (lalaly):

is that what you need?

OpenStudy (anonymous):

Whatt Ithought It Was y=2x+1

OpenStudy (lalaly):

\[5x+10=10y\]\[10y=5x+10\]divide all by 10 \[\frac{10}{10}y=\frac{5}{10}x+\frac{5}{10}\] \[y=\frac{1}{2}x+1\]

OpenStudy (lalaly):

+1/2 in the endnot 1 sory